# A new perspective on the equivalence between Weak and Strong Spatial Mixing in two dimensions

### v1

@@ 4 Application II: FK percolation
One checks the Hypotheses (Mix1)(Mix2)(Mar1)(Mar2)(Mar3) separately in the strongly subcritical and strongly supercritical cases. Then, application of Theorem 1.1 to the collection of measure $\Psi_{\Lambda}^{\pi}$ with $\pi\in\{0,1\}$ defined in the previous subsection gives the result as in the proof of Theorem 3.1. Let $l>4R$ ($R$ the range of the percolation model) be a large enough number.
@@
Following Remark 1.1, we only need to define the sets $\mathrm{Ma}_{u}$, and the spin values $\mathrm{a}_{u},\mathrm{a}_{u,\Lambda}^{t}$, $u\in\mathbb{L}_{l}$. Let $\mathrm{Ma}_{u}$ be the event that the configuration $\eta_{\Lambda_{3l+1}(u)}$ has no connections from $\Lambda_{l}(u)$ to $\Lambda_{3l/2}(u)^{c}$. The elements $\mathrm{a}_{u},\mathrm{a}_{u,\Lambda}^{t}$ are then constant 00 configurations. Finite energy gives the wanted properties of $\mathrm{a}_{u},\mathrm{a}_{u,\Lambda}^{t}$, and strong sub-criticality gives the wanted large probability of $\mathrm{Ma}_{u}$ once $l$ is taken large enough.
Hypotheses (Mix1)(Mix2) are then a direct consequence of Theorem 1.2 and weak mixing.
@@ 5 Application III: Hard core models
As in Section 3, take $l$ strictly bigger than the diameter of $D$. Then, (Mar1)(Mar2) hold with $p_{\mathrm{bulk}}=1$. The only change with respect to Section 3 is that one takes $\mathrm{a}_{u}=0$ (constant 00 configuration on $\Lambda_{l}(u)$) and $\mathrm{a}_{u,\Lambda}^{\xi}=0$ (constant 00 configuration on $\Lambda_{l}(u)\cap\Lambda$). Following the same procedure as in Section 3, one obtains that weak mixing (satisfying (11)) implies strong mixing (satisfying (12)) for volumes that are large enough squares.
@@ Appendix A Bernoulli percolation
The results/proofs presented here are mostly exercises. They are included for completeness.
@@ A.2 Percolation estimates
Let $G=(V,E)$ be a connected graph. Let $P_{p}$ be the site Bernoulli percolation measure of parameter $p$ on $G$. Let $X\sim P_{p}$. Then, for any $\lambda\geq$ 00, and any $A\subset V$,
@@
$A=\{\text{there is exactly one connected component of size }\geq n/4\},$
@@
Not that the closed sites form a percolation $1-p$ which is as close to 00 as wanted. First, the event $A^{c}$ implies that at least one vertex belongs to a closed *-cluster of size at least $n/4$. Applying Lemma A.3 with $d=8$, $\lambda=13$, and using Chebychev’s inequality, one obtains that for any $p$ close enough to $1$, and $n$ large enough,
@@ Appendix B Conditional Factorization to Ratio Mixing
Let $\Omega_{i},\,i=1,2$ be finite sets. Let $\Omega=\Omega_{1}\times\Omega_{2}$ and $\mathcal{F}_{i}=\{A\subset\Omega_{i}\}$, =$\mathcal{F}=\{A\subset\Omega\}$. Let $\mu,\nu$ probability measures on $(\Omega,\mathcal{F})$. Let

### v2

@@ 4 Application II: FK percolation
One checks the Hypotheses (Mix1), (Mix2), (Mar1), (Mar2), (Mar3) separately in the strongly subcritical and strongly supercritical cases. Then, application of Theorem 1.1 to the collection of measure $\Psi_{\Lambda}^{\pi}$ with $\pi\in\{0,1\}$ defined in the previous subsection gives the result as in the proof of Theorem 3.1. Let $l>4R$ ($R$ the range of the percolation model) be a large enough number.
@@
Following Remark 1.1, we only need to define the sets $\mathrm{Ma}_{u}$, and the spin values $\mathrm{a}_{u},\mathrm{a}_{u,\Lambda}^{t}$, $u\in\mathbb{L}_{l}$. Let $\mathrm{Ma}_{u}$ be the event that the configuration $\eta_{\Lambda_{3l+1}(u)}$ has no connections from $\Lambda_{l}(u)$ to $\Lambda_{3l/2}(u)^{c}$. The elements $\mathrm{a}_{u},\mathrm{a}_{u,\Lambda}^{t}$ are then constant 0 configurations. Finite energy gives the wanted properties of $\mathrm{a}_{u},\mathrm{a}_{u,\Lambda}^{t}$, and strong sub-criticality gives the wanted large probability of $\mathrm{Ma}_{u}$ once $l$ is taken large enough.
Hypotheses (Mix1), (Mix2) are then a direct consequence of Theorem 1.2 and weak mixing.
@@ 5 Application III: Hard core models
As in Section 3, take $l$ strictly bigger than the diameter of $D$. Then, (Mar1), (Mar2) hold with $p_{\mathrm{bulk}}=1$. The only change with respect to Section 3 is that one takes $\mathrm{a}_{u}=0$ (constant 0 configuration on $\Lambda_{l}(u)$) and $\mathrm{a}_{u,\Lambda}^{\xi}=0$ (constant 0 configuration on $\Lambda_{l}(u)\cap\Lambda$). Following the same procedure as in Section 3, one obtains that weak mixing (satisfying (12)) implies strong mixing (satisfying (13)) for volumes that are large enough squares.
@@ Appendix A Bernoulli percolation
The results/proofs presented here are probably exercises (or even trivialities!) for percolation experts. They are included for completeness.
@@ A.2 Percolation estimates
Let $G=(V,E)$ be a connected graph. Let $P_{p}$ be the site Bernoulli percolation measure of parameter $p$ on $G$. Let $X\sim P_{p}$. Then, for any $\lambda\geq$ 0, and any $A\subset V$,
@@
[…]
@@
Not that the closed sites form a percolation 1p1-p which is as close to 0 as wanted. First, the event $A^{c}$ implies that at least one vertex belongs to a closed *-cluster of size at least $n/4$. Applying Lemma A.3 with $d=8$, $\lambda=13$, and using Chebychev’s inequality, one obtains that for any $p$ close enough to $1$, and $n$ large enough,
@@
A.3 Boundary inhomogeneous percolation
This section contains a proof of exponential decay for (very) subcritical site Bernoulli percolation with inhomogeneities at the boundary of a square box. This is the result used in the Applications.
Lemma A.5.
Let $p<0.001$, and $q\in[0,1)$. Then, there are $c>0,C\geq 0$ such that for any $n\geq 1$, and any $x,y\in\Lambda_{n}\equiv\{-n,\dots,n\}^{2}$,
$P_{n;p,q}(x\leftrightarrow_{*}y)\leq Ce^{-c|x-y|}$
where $P_{n;p,q}$ is the law of a site Bernoulli percolation on $\Lambda_{n}$ with parameters $p$ in $\Lambda_{n}\setminus\partial^{\mathrm{i}}\Lambda_{n}$, and $q$ in $\partial^{\mathrm{i}}\Lambda_{n}$.
Proof.
Let $X\sim P_{n;p,q}$. Let $\mathring{X}$ be the restriction of $X$ to $\Lambda_{n}\setminus\partial^{\mathrm{i}}\Lambda_{n}$ (a site percolation Bernoulli with parameter $p$). As $p<0.001$, one has that there is $c>0$ such that for any $u,v\in\Lambda_{n}\setminus\partial^{\mathrm{i}}\Lambda_{n}$,
$P\big(u\xleftrightarrow{\mathring{X}}_{*}v\big)\leq e^{-c|u-v|}.$
In particular, for any $x,y\in\Lambda_{n}$,
$P\big(x\xleftrightarrow{X}_{*}y\big)$ $\leq P\big(x\xleftrightarrow{\mathring{X}}_{*}y\big)+\sum_{u,v\in\partial^{\mathrm{i}}\Lambda_{n}}P\big(\{x\xleftrightarrow{\mathring{X}}_{*}\sim u\}\circ\{u\xleftrightarrow{X}_{*}v\}\circ\{y\xleftrightarrow{\mathring{X}}_{*}\sim v\}\big)$
$\leq e^{-c|x-y|}+\sum_{u,v\in\partial^{\mathrm{i}}\Lambda_{n}}P\big(x\xleftrightarrow{\mathring{X}}_{*}\sim u\big)P\big(u\xleftrightarrow{X}_{*}v\big)P\big(y\xleftrightarrow{\mathring{X}}_{*}\sim v\big)$
$\leq e^{-c|x-y|}+9\sum_{u,v\in\partial^{\mathrm{i}}\Lambda_{n}}e^{-c|x-u|}P\big(u\xleftrightarrow{X}_{*}v\big)e^{-c|y-v|}$
where $x\leftrightarrow_{*}\sim y$ is the event that there is a *-connection from $x$ to one of the *-neighbours of $y$, $\circ$ denotes disjoint occurrence, we used the BK inequality in the second inequality, and a union bound over the neighbours of $u,v$ that are in $\Lambda_{n}\setminus\partial^{\mathrm{i}}\Lambda_{n}$ in the last. To get the claim, it is sufficient to show that there is $c>0$ such that for any $x,y\in\partial^{\mathrm{i}}\Lambda_{n}$,
$P\big(x\xleftrightarrow{X}_{*}y\big)\leq e^{-c|x-y|}.$
For $v\in\partial^{\mathrm{i}}\Lambda_{n}$, let
$\mathcal{C}_{v}=\{x\in\Lambda_{n}\setminus\partial^{\mathrm{i}}\Lambda_{n}:\ \exists y\in\Lambda_{n}\setminus\partial^{\mathrm{i}}\Lambda_{n},\|y-v\|_{\infty}=1,x\xleftrightarrow{\mathring{X}}_{*}y\}.$
$\mathcal{C}_{v}$ is measurable in terms of $\mathring{X}$ and is contained in $\Lambda_{n}\setminus\partial^{\mathrm{i}}\Lambda_{n}$. Define
$\mathcal{C}_{\partial}=\cup_{v\in\partial^{\mathrm{i}}\Lambda_{n}}\mathcal{C}_{v}.$
Let now $x,y\in\partial^{\mathrm{i}}\Lambda_{n}$. Call $\{u,v\}\subset\partial^{\mathrm{i}}\Lambda_{n}$ a $(x,y)$-pre-cut if $x,y$ lie in different *-connected components of $(\partial^{\mathrm{i}}\Lambda_{n}\setminus\{u,v\})\cup\mathcal{C}_{\partial}$, see Figure 9.
Figure 9: Orange disks: $\partial^{\mathrm{i}}\Lambda_{n}$, blue disks: $\mathcal{C}_{\partial}$, black squares: a $(x,y)$-pre-cut.
Observe that conditionally on the existence of at least $k$ disjoint $(x,y)$-pre-cuts, the probability of $x\leftrightarrow_{*}y$ is at most $(1-(1-q)^{2})^{k}$ as at least one site in each of them has to be open in $X$ to realize the connection. The wanted result is then reduced to showing that there is at least $\rho|x-y|$ $(x,y)$-pre-cuts with probability at least $1-e^{-c|x-y|}$ for some $c,\rho>0$. Let $Y_{v}$, $v\in\partial^{\mathrm{i}}\Lambda_{n}$, be an i.i.d. collection of $\mathbb{Z}$-valued random variables with
$P(Y_{v}=k)=P_{p}\big(X_{v}=1,\max\{\|u\|_{\infty}:\ u\leftrightarrow_{*}0\}=k\big)\leq e^{-ck},\ k\geq 0,$
$P(Y_{v}=-1)=P_{p}(X_{v}=0),$
where $P_{p}$ is a Bernoulli site percolation of parameter $p$ on $\mathbb{Z}^{2}$, and $c>0$ (this last fact is due to $p<0.001$). Define
$\mathcal{D}=\bigcup_{v\in\partial^{\mathrm{i}}\Lambda_{n}}(v+\Lambda_{Y_{v}})$
where $\Lambda_{-1}=\varnothing$. This set stochastically dominates $\mathcal{C}_{\partial}$. In particular, the wanted bound on the presence of $(x,y)$-pre-cuts is implied by showing that on each of the two paths going from $x$ to $y$ in $\partial^{\mathrm{i}}\Lambda_{n}$, there are at least $\rho|x-y|$ points that are not in $\mathcal{D}$ with probability at least $1-e^{-c|x-y|}$ for some $c,\rho>0$. But this last estimate is a simple one-dimensional percolation estimate for long range stick percolation with exponential tails. The reader in need is invited to look at the Appendix of [13] for details on how to implement this kind of argument. ∎
Remark A.1.
The above argument can be improved up to $p$ smaller than the critical point for *-percolation. The choice $0.001$ is arbitrary and sufficient to implement a Peierls argument to obtain exponential bounds for the point-to-boundary percolation probabilities.
@@ Appendix B Conditional Factorization to Ratio Mixing
Let $\Omega_{i},\,i=1,2$ be finite sets. Let $\Omega=\Omega_{1}\times\Omega_{2}$ and $\mathcal{F}_{i}=\{A\subset\Omega_{i}\}$, $\mathcal{F}=\{A\subset\Omega\}$. Let $\mu,\nu$ probability measures on $(\Omega,\mathcal{F})$. Let